Show all work please.1. U(x, y) x,ax,1-a) a. Solve for the marshallian demands for x, and x2, as functions of p1, p2, and m. (Hint: your solutions
Answer #1

(a) UCXI,X2) = x,a xal-a ov - Mo, ax4-1 x, -а OX1 ao - mor= (1-2) X, 9 x2 t-a-1 = (1-2) X, 9 xz a MRS = - мо, MU2 ax,- х-4 (Pix lat la) zm 1 = m Pixn= am xi² am *2= line you lanmou) X2 = (-a)m P2 X = am 2x % own price elasticity of ad d o ExiPi = 0xincome elasticity of demand Ex, m = em = 1 c) a=0.2, m=100, P1=2, P2 = 8 xi= am - x = A = (220 b) = lo (0.2) (150) Po x2 = (-